# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

### v2

@@ 5.1.2 Benchmark case 2: The GC Model
Figure 3: Comparison of phase-field (PF) simulation results with the sharp interface (SI) results via FEM method for the GC case. Figures (a) and (b) showcase the variations in grain boundary potential, dopant concentration, and oxygen vacancy concentration in the core as functions of $\Delta g_{\text{dop}}$. Figures (c), (d), and (e) detail the spatial distributions of dopant and oxygen vacancy concentrations in the bulk phase and the electrostatic potential. Figure (f) depicts the area density of net charge in the core, also as a function of $\Delta g_{\text{dop}}$. $\Delta g_{\text{V}_{\text{O}}^{..}}$ was kept at -1.0 eV for all calculations.
@@ 5.2 GB space charge layers at quasi-equilibrium
In this section, we investigate systematically the solute drag effect caused by an asymmetric SCL at a quasi-statically moving GB in a STO bicrystal. The solute drag phenomenon arises when point defects, segregated within a moving core in a quasi-equilibrium state, lead to additional free energy dissipation to maintain motion. At low core velocities, these defects can keep pace with the core, significantly increasing the necessary driving force. However, beyond a critical velocity, the defects lag behind, remaining in the bulk phases and causing the breakdown of symmetric concentration and electrostatic potential profiles across the SCL.
To investigate the formation of asymmetric SCL due to solute drag effects under quasi-equilibrium conditions, a constant core velocity (${v}_{\text{c}}$) along the simulation domain is assigned as a parameter in the simulations. The velocity of the core ranges from 0.01 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$ to study its influence on the SCL. The defect chemistry parameters $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ in this case are $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}$ $\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$, respectively. Simulations are conducted at 1623 K, with a relative permittivity ($\epsilon_{\text{r}}$) of STO set to 56, based on [38]. The diffusivities of oxygen vacancies and dopant in the bulk phase are specified as $1.35\times 10^{-9}$ $\text{m}^{2}\text{s}^{-1}$ and $1.3\times 10^{-11}$ $\text{m}^{2}\text{s}^{-1}$, respectively [38]. The simulation domain and boundary conditions remain consistent with the previous case.
In alignment with Vikrant [38], the total driving force $F_{\text{T}}$ for the core motion is given by
$F_{\text{T}}=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{V}_{\text{O}}^{..}}}\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{dop}}}\frac{\partial C_{\text{dop}}}{\partial x}\text{d}x.$ (60)
The second and third terms in Eq. 60 denote as the electrochemical drag force, which accounts for the electrostatic and chemical contributions. The integration in Eq. 60 are solved numerically.
In Fig. 4, we illustrate the concentration profiles of oxygen vacancies and dopant, as well as the charge density and electrostatic potential as functions of $x$. From Fig. 4(a) to Fig. 4(e), we progressively increase the velocity of the core from 0 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$. The direction of the core motion is from left to right along $x$-axis. In the following context, the left side of the core is named as the growing grain, while the right side is called the shrinking grain.
In Fig. 4(a), the symmetric SCL formation is presented at $v_{c}=0$ $\text{nm}\,\text{s}^{-1}$. However, the symmetric pattern begins to deteriorate even at very low velocity, e.g. 0.01 $\text{nm}\,\text{s}^{-1}$, as shown in Fig. 4(b1), (b2) and (b3). Compared with the concentration profiles at an equilibrium state, the accumulation zone of dopant in the growing grain is reduced, while that of the shrinking grain increases. This asymmetry arises, since during the core motion, dopant cations in the shrinking grain are captured by core due to their low diffusivity. Driving the motion of extra dopant cations within the core demands extra free energy dissipation. Therefore, the dopant cations flowing from the core region into the growing grain lead to the enhancement of the accumulation zone in the growing grain. However, the concentration of dopant in the core decreases from $7.8\times 10^{25}$ $\text{m}^{-3}$ ($v_{c}=0$ in Fig. 4(a1)) to $6.8\times 10^{28}$ $\text{m}^{-3}$ at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$. In the growing grain far from the core, dopant concentration reaches $7.2\times 10^{25}$ $\text{m}^{-3}$, which is higher than equilibrium concentration, i.e. $c_{\text{dop},\text{b}}(\infty)$, as shown in Fig. 4(a1). In Fig. 4(b2), we plot the calculated charge density $\rho=e(z_{\text{dop}}c_{\text{dop}}+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..}})$. The charge density in the core decreases. Compared with the equilibrium case, we also observe the formation of less (more) negative charged zone in the shrinking (growing) grain, respectively. In Fig. 4(b3), the grain boundary potential increases from 0.248 V ($v_{c}=0$ $\text{nm}\,\text{s}^{-1}$) to 0.265 V at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$. Moreover, an electrostatic potential different of 0.011 V develops between the shrinking and growing grains.
When we increase the velocity of the core to 0.5 $\text{nm}\,\text{s}^{-1}$, dopant cations are not capable of catching the movement of core, as depicted in Fig. 4(c1). Consequently, the dopant concentration within the core decreases substantially, resulting in a smaller depletion zone ahead of the core and a reduced accumulation zone behind it. Similarly, the concentration of oxygen vacancies within the core also decreases significantly. Correspondingly, as shown in Fig. 4(c2), the charge density in the core drops to $8.6\times 10^{-7}$ $\text{C}\text{m}^{-3}$, which contrasts with the increase in charge density in the core observed in Virkant’s work [38] at intermediate velocities. We attribute the difference to the defect chemistry incorporated in the present phase field model.
When the core velocity increases to 2 $\text{nm}\,\text{s}^{-1}$, the segregation of dopant in the core becomes negligible, as shown in Fig. 4(d1). The concentration of dopant on both the left and right sides of the bulk phase returns to the equilibrium value, i.e. $c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$. In Fig. 4(d2), due to the continuous decrease in charge density within the core, the negatively charged zone is diminished. Furthermore, the grain boundary potential increases to 0.446 V and the electrostatic potential difference between the left and right sides reaches 0.122 V.
At a high core velocity, such as 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration profile becomes completely flat, while the oxygen vacancy concentration profile exhibits strong asymmetry, as depicted in Fig. 4(e1). The concentration of oxygen vacancies within the core decreases significantly to $2.46\times 10^{26}$ $\text{m}^{-3}$. In the shrinking grain, the depletion zone of oxygen vacancies remains sharp, about 2.5 nm, but it extends to approximately 20 nm in the growing grain. Consequently, as shown in Fig. 4(e2), the negatively charged zone in the shrinking grain is much smaller, and the negative charge region in the growing grain is negligible. In Fig. 4(e3), the electrostatic potential is illustrated. The grain boundary potential increases to 1.02 V and decays very slowly in the growing grain. In addition, the electrostatic potential difference becomes more pronounced, reaching 0.995 V.
In summary, at small core velocity, the asymmetric formation of SCL is mainly caused by dopant coming apart from the core. The charge density in the negatively charged SCL and in the positively charged core both decrease. As a result, the electrostatic potential in the shrinking grain increases to a positive value. At high core velocity, dopant cations do not segregate to the core. The oxygen vacancy concentration profile is strongly asymmetric. This asymmetry causes an extended diffusion layer in the growing grain and elevates the electrostatic potentials in both the core and the shrinking grain.
@@
Figure 4: Asymmetric SCL formation in a STO bicrystal system with $c_{\text{dop},\text{b}}(\infty)=7.0\times 10^{25}$ $\text{m}^{-3}$, $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV and $\Delta g_{\text{dop}}=-0.5$ eV at quasi-equilibrium state under different constant core velocity (a) $v_{\text{c}}=0$ $\text{nm}\,\text{s}^{-1}$, (b) $v_{\text{c}}=0.01$ $\text{nm}\,\text{s}^{-1}$, (c) $v_{\text{c}}=0.5$ $\text{nm}\,\text{s}^{-1}$, (d) $v_{\text{c}}=2$ $\text{nm}\,\text{s}^{-1}$, (e) $v_{\text{c}}=100$ $\text{nm}\,\text{s}^{-1}$. Each row has three sub-figures, depicting the distribution of dopant and oxygen vacancy concentrations, total charge density and electrostatic potential, respectively.
@@
In order to further investigate solute drag effects and identify the critical core velocity at which a velocity jump occurs during sintering process, we present the core velocity as a function of the total driving force calculated via Eq. 60 in Fig. 5. The segregation energy difference, $\Delta g_{\text{dop}}$ and $\Delta g_{\text{V}_{\text{O}}^{..}}$, allow direct tuning of the concentrations of oxygen vacancy and dopant in the core, thereby impacting the critical total driving force required for a specific core velocity. In Fig. 5(a), we hold $\Delta g_{\text{dop}}$ and $c_{\text{dop},\text{b}}(\infty)$ constant at -0.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and progressively decrease $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -0.5 eV to -1.75 eV. During the grain shrinking process, the driving force resulting from the GB curvature becomes increasingly significant. Consequently, the velocity of the GB core evolves from a low value to a high value. At low core velocities, a more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$ induces higher concentration of oxygen vacancy in the core consequently demanding a larger driving force to maintain the same velocity. As the core velocity increases to 1 $\text{nm}\,\text{s}^{-1}$, the dopant cations are unable to keep pace with the core movement, resulting in a nearly flat dopant concentration within the core. Hence, this leads to the same linear relationship between the total driving force and the core velocity for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ at high velocity. On the other hand, for fixed $\Delta g_{\text{V}_{\text{O}}^{..}}$, such as $\Delta g_{\text{V}_{\text{O}}^{..}}$=-1.75 eV, we clearly observe the velocity jumps from 0.065 $\text{nm}\,\text{s}^{-1}$ to approximate 1.5 $\text{nm}\,\text{s}^{-1}$ at $1.45\times 10^{7}$ $\text{N}\text{m}^{-2}$. As we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ to -1.0 eV, we hardly observe the velocity jump. When $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -0.5 eV, both the concentrations of dopant and oxygen vacancy in the core decrease significantly. The solute drag effect therefore becomes negligible. Fig. 5(b) shows the influence of $\Delta g_{\text{dop}}$ on solute drag effects when $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ are held constant at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$. In contrast to Fig. 5(a), we can evidently observe velocity jumps for all $\Delta g_{\text{dop}}$. As we decrease $\Delta g_{\text{dop}}$ to less negative value, such as from -0.2 eV to -0.3 eV, the enhancement of the solute drag effect is not significant. In Fig. 5(c), we explore the impact of the dopant concentration in the bulk far from the core, denoted as $c_{\text{dop},\text{b}}(\infty)$. We fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ at -1.5 eV and -0.5 eV. As $c_{\text{dop},\text{b}}(\infty)$ increases, a larger driving force is necessary to counterbalance the solute drag effect to achieve the same core velocity. Additionally, a lower critical velocity is observed when the velocity jump occurs at smaller values of $c_{\text{dop},\text{b}}(\infty)$. Specifically, the critical velocity increases from 0.065 $\text{nm}\,\text{s}^{-1}$ to 0.095 $\text{nm}\,\text{s}^{-1}$, as $c_{\text{dop},\text{b}}(\infty)$ is raised from $4\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$.
@@
In the sintering process, the driving forces are initially large and decrease as the grain size increases. In other words, the GB core velocity starts on the upper right side of the S-shaped curves in Fig. 5 at the initial stage of the grain growth. As the grain size increases, the GB core velocity decreases to the critical velocity. At this point, a transition from high velocity to low velocity occurs at the critical value. In Fig. 5(a), we observe GB core velocity jumps when $\Delta g_{\text{V}_{\text{O}}^{..}}=$1.5 eV and -1.75 eV. The critical velocities for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ are approximately 0.35 $\text{nm}\,\text{s}^{-1}$. Additionally, the GB core velocity jumps to a lower value when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is more negative. In Fig. 5(b), the critical GB core velocity increases from 0.15 $\text{nm}\,\text{s}^{-1}$ to 0.65 $\text{nm}\,\text{s}^{-1}$ as $\Delta g_{\text{dop}}$ decreases from -0.3 eV to -0.7 eV. In Fig. 5(c), a larger $c_{\text{dop},\text{b}}(\infty)$ leads to a higher critical velocity.
@@
5.3 Abnormal grain growth

### v3

@@ 5.1.2 Benchmark case 2: GC Model
Figure 3: Comparison of phase-field (PF) simulation results with the abrupt GB||SCL sharp interface (SI) results via FEM method for the GC case. Figures (a), (b) and (c) showcase the variations in grain boundary potential, dopant concentration, and oxygen vacancy concentration in the core as functions of $\Delta g_{\text{dop}}$. Figures (d), (e), and (f) detail the spatial distributions of dopant and oxygen vacancy concentrations in the bulk phase and the electrostatic potential. $\Delta g_{\text{V}_{\text{O}}^{..}}$ was kept at -1.5 eV for all calculations.
@@ 5.2 GB space charge layers at quasi-equilibrium
In this section, we systematically investigate the solute drag effect induced by an asymmetric SCL at a quasi-statically moving GB in a STO bicrystal case. This analysis is conducted using the defect-chemistry-consistent phase-field model, which has been validated against sharp interface predictions in previous section. While the impacts of GB misorientation have been extensively studied in Ref. vikrant2020electrochemical , our focus here is on the influence of defect chemistry parameters, such as the segregation energies of oxygen vacancy and dopant, on SCL formation and solute drag effects at the quasi-equilibrium state in a bicrystal case.
The phase-field simulation parameters to reproduce the SCLs at quasi-equilibrium state are listed in Table 4. For FEM implementation, the simulation domain is a 1D space extending from x=780x=-780 nm to $x=780$ nm, with the GB core initially positioned at x=390x=-390 nm ($t=0$). The domain size is chosen to be sufficiently large to ensure the movement of SCL is not hindered by the boundary of simulation box. The domain is discretized into 20,000 elements, each with a uniform size of 0.078 nm. For the electrostatic potential, a Neumann boundary condition is applied at the right boundary, while the left boundary is grounded. Similarly, Neumann boundary conditions are imposed on the concentration field at both boundaries. Each simulation is performed on a single CPU node with 16 cores and 8 GB of memory, requiring approximately 12 hours to reach the quasi-equilibrium state.
Table 4: The defect chemistry parameters for the 1D phase-field simulations to reproduce quasi-equilibrium SCLs.
Parameters Value
TT/K 1623
$D_{\text{V}_{\text{O}}^{..}}$$\text{m}^{-2}\text{s}^{-1}$ $1.35\times 10^{-9}$
$D_{\text{dop}}$$\text{m}^{-2}\text{s}^{-1}$ $1.3\times 10^{-11}$
aa/nm 0.39
$\omega_{c}$/nm 0.78
$\epsilon_{r}$ 56
$c_{\text{dop},\text{b}}(\infty)$/ $\text{m}^{-3}$ $2\times 10^{25}$ $\sim$ $2\times 10^{26}$
$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$/ $\text{m}^{-3}$ $0.5\times c_{\text{dop},\text{b}}(\infty)$
$\Delta g_{\text{V}_{\text{O}}^{..}}$/eV -0.5 $\sim$ -1.75
$\Delta g_{\text{dop}}$/eV -0.2 $\sim$ -0.7
$N_{\text{V}_{\text{O}}^{..},\text{b}}$/ $\text{m}^{-3}$ $5.1\times 10^{28}$
$N_{\text{dop},\text{b}}$/ $\text{m}^{-3}$ $1.68\times 10^{28}$
$N_{\text{V}_{\text{O}}^{..},\text{c}}$/ $\text{m}^{-3}$ $1\times 10^{27}$
$N_{\text{dop},\text{c}}$/ $\text{m}^{-3}$ $1.68\times 10^{27}$
To investigate the formation of asymmetric SCLs induced by solute drag effects under quasi-equilibrium conditions, a constant core velocity (${v}_{\text{c}}$) is prescribed as a parameter in the simulations. The core velocity ranges from 0.01 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$ to study its influence on the SCL. The simulation results, shown in Fig. 4, illustrate the distributions of acceptor dopant concentration, oxygen vacancy concentration, charge density, and electrostatic potential as functions of $x$. The charge density is calculated as $\rho(x)=e(z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}c_{\text{dop}})$. The GB core moves from left to right along the $x$-axis. The acceptor dopant plays a crucial role in solute drag effects due to its significantly lower diffusivity compared to oxygen vacancy. Adjusting the segregation energy of the dopant significantly alters its concentration distribution. To examine the impact of segregation behavior, two segregation energies for the acceptor dopant, 0 eV and 0.5-0.5 eV, are selected to generate different SCLs. The segregation energy of oxygen vacancy is fixed at 1.5-1.5 eV. In the following discussion, the region to the left of the GB core is referred to as the growing grain, while the region to the right is referred to as the shrinking grain.
The SCLs at the equilibrium state (i.e., $v_{\text{c}}$ = 0 $\text{nm}\,\text{s}^{-1}$) are shown in Fig. 4(a) as a reference. The two different segregation energies for the acceptor dopant lead to distinct concentration distributions, as illustrated in Fig. 4(a1). For $\Delta g_{\text{dop}}=0$ eV, the acceptor dopant concentration exhibits a depression at the GB core, while $\Delta g_{\text{dop}}=-0.5$ eV results in segregation. Moreover, the charge density in the GB core is higher for $\Delta g_{\text{dop}}=0$ eV than for $\Delta g_{\text{dop}}=-0.5$ eV. Consequently, the GB potential for $\Delta g_{\text{dop}}=-0.5$ eV is smaller than that for $\Delta g_{\text{dop}}=0$ eV.
Compared with symmetric SCLs, a moving GB core introduces notable differences. First, the concentration profiles of the dopant are highly sensitive to GB velocity. The symmetric pattern begins to deteriorate even at very low velocities, such as 0.01 $\text{nm}\,\text{s}^{-1}$, as shown in Fig. 4(b1). As a result, the dopant accumulation zone increases in the growing grain but decreases in the shrinking grain. At higher core velocities, dopant cations fail to keep up with the movement of the GB core, leading to a significant decrease in dopant concentration within the core, as illustrated in Fig. 4(c1), (d1), and (e1). When the core velocity reaches $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration profile becomes nearly flat, as depicted in Fig. 4(e1). Second, significant asymmetry in the oxygen vacancy concentration arises only at very high GB velocities, due to its high diffusivity. As depicted in Fig. 4(b2), (c2), and (d2), the oxygen vacancy concentration profiles exhibit only slight asymmetry at lower velocities. However, when $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the concentration of oxygen vacancy in the depletion zone of the growing grain increases notably, as shown in Fig. 4 (e2). Simultaneously, the depletion zone in the growing grain becomes substantially elongated, whereas it contracts in the shrinking grain. Third, as the GB velocity increases, not only does the GB potential rise, but a difference in electrostatic potential between the growing and shrinking grains also emerges, as shown in Fig. 4(b4)–(e4). Additionally, the decay length of the electrostatic potential in the growing grain becomes significantly longer when GB velocity increases, resulting in an elongated SCL length in the growing grain.
The segregation energy of the acceptor dopant plays a crucial role in SCL formation when the GB core is moving. Different values of $\Delta g_{\text{dop}}$, namely 0.5-0.5 eV and 0 eV, lead to distinct dopant concentration distributions. In Fig. 4(b1), when the GB velocity $v_{\text{c}}$ is $0.01$ $\text{nm}\,\text{s}^{-1}$, the accumulation zones in the SCL for both segregation energies exhibit the same trend. However, as $v_{\text{c}}$ increases to $0.5$ $\text{nm}\,\text{s}^{-1}$ and 2 $\text{nm}\,\text{s}^{-1}$, notable differences emerge as shown in Fig. 4(c1) and (d1). For $\Delta g_{\text{dop}}=-0.5$ eV, the accumulation zone in the shrinking grain transforms into a depletion zone, where the dopant concentration becomes lower than in the bulk. Meanwhile, the accumulation zone in the growing grain persists. In contrast, for $\Delta g_{\text{dop}}=0$ eV, the opposite behavior is observed. The accumulation zone remains in the shrinking grain, while in the growing grain, it diminishes into a depletion zone. When $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration becomes flat for both cases. The influence of the segregation energy of the acceptor dopant on the electrostatic potential is significant only at low GB velocities. A more negative $\Delta g_{\text{dop}}$ leads to a smaller GB potential. At high GB velocities, where the dopant concentration profiles flatten, their effect on the electrostatic potential distribution diminishes, as shown in Fig. 4(c4), (d4), and (e4).
@@
Figure 4: Asymmetric SCL formation in a STO bicrystal system with $c_{\text{dop},\text{b}}(\infty)=7.0\times 10^{25}$ $\text{m}^{-3}$, $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV and two $\Delta g_{\text{dop}}=$ 0 and -0.5 eV at quasi-equilibrium state under different constant core velocity (a) $v_{\text{c}}=0$ $\text{nm}\,\text{s}^{-1}$, (b) $v_{\text{c}}=0.01$ $\text{nm}\,\text{s}^{-1}$, (c) $v_{\text{c}}=0.5$ $\text{nm}\,\text{s}^{-1}$, (d) $v_{\text{c}}=2$ $\text{nm}\,\text{s}^{-1}$, (e) $v_{\text{c}}=100$ $\text{nm}\,\text{s}^{-1}$. Each row has four sub-figures, depicting the distribution of dopant concentration, oxygen vacancy concentration, total charge density and electrostatic potential, respectively. In this figure, blue line indicates $\Delta g_{\text{dop}}$ = -0.5 eV, while red line is $\Delta g_{\text{dop}}$ = 0 eV.
The influence of defect chemistry parameters, especially segregation energy of dopant, on the formation behavior of different SCLs has been demonstrated. In order to further investigate solute drag effects and identify the critical core velocity at which a velocity jump occurs during sintering process, we present the core velocity as a function of the total driving force calculated via the following equation taken from Ref. vikrant2020electrochemical
$\begin{split}F_{\text{T}}&=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{V}_{\text{O}}^{..}}}\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{dop}}}\frac{\partial C_{\text{dop}}}{\partial x}\text{d}x\\ &=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x-\int_{-\infty}^{\infty}\left[C_{\text{V}_{\text{O}}^{..}}-C_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\right]\frac{\partial\mu_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x-\int_{-\infty}^{\infty}\left[C_{\text{dop}}-C_{\text{dop},\text{b}}(\infty)\right]\frac{\partial\mu_{\text{dop}}}{\partial x}\text{d}x.\end{split}$ (65)
The second and third terms in Eq. 65 denote as the electrochemical drag force, which accounts for the electrostatic and chemical contributions when the expressions of electrochemical potential is employed. The integration in Eq. 65 are solved numerically.
@@
The segregation energy difference, $\Delta g_{\text{dop}}$ and $\Delta g_{\text{V}_{\text{O}}^{..}}$, allow direct tuning of the concentrations of oxygen vacancy and dopant in the core, thereby impacting the critical total driving force required for a specific core velocity. In Fig. 5(a), we hold $\Delta g_{\text{dop}}$ and $c_{\text{dop},\text{b}}(\infty)$ constant at -0.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and progressively decrease $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -0.5 eV to -1.75 eV. During the grain shrinking process, the driving force resulting from the GB curvature becomes increasingly significant. Consequently, the velocity of the GB core evolves from a low value to a high value. At low core velocities, a more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$ induces higher concentration of oxygen vacancy in the core consequently demanding a larger driving force to maintain the same velocity. As the core velocity increases to 1 $\text{nm}\,\text{s}^{-1}$, the dopant cations are unable to keep pace with the core movement, resulting in a nearly flat dopant concentration within the core. Hence, this leads to the same linear relationship between the total driving force and the core velocity for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ at high velocity. On the other hand, for fixed $\Delta g_{\text{V}_{\text{O}}^{..}}$, such as $\Delta g_{\text{V}_{\text{O}}^{..}}$=-1.75 eV, we clearly observe the velocity jumps from 0.065 $\text{nm}\,\text{s}^{-1}$ to approximate 1.5 $\text{nm}\,\text{s}^{-1}$ at $1.45\times 10^{7}$ $\text{N}\text{m}^{-2}$. As we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ to -1.0 eV, we hardly observe the velocity jump. When $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -0.5 eV, both the concentrations of dopant and oxygen vacancy in the core decrease significantly. The solute drag effect therefore becomes negligible. Therefore, decreasing the segregation energy of oxygen vacancy could be an effective way to eliminate the solute drag effects. Fig. 5(b) shows the influence of $\Delta g_{\text{dop}}$ on solute drag effects when $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ are held constant at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$. In contrast to Fig. 5(a), we can evidently observe velocity jumps for all $\Delta g_{\text{dop}}$. When the segregation energy of dopant decrease to -0.2 eV, we can still observe the solute drag effects. In Fig. 5(c), we explore the impact of the dopant concentration in the bulk far from the core, denoted as $c_{\text{dop},\text{b}}(\infty)$. We fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ at -1.5 eV and -0.5 eV. We clearly find the solute drag effect is strongly dependent on dopant concentration as evidenced in Ref. zahler2023grain . When $c_{\text{dop},\text{b}}=2\times 10^{26}\text{m}^{-3}$ (corresponding to 0.12% site fraction), solute drag effect still remains. As $c_{\text{dop},\text{b}}(\infty)$ increases, a larger driving force is necessary to counterbalance the solute drag effect to achieve the same core velocity. Additionally, a lower critical velocity is observed when the velocity jump occurs at smaller values of $c_{\text{dop},\text{b}}(\infty)$. Specifically, the critical velocity increases from 0.065 $\text{nm}\,\text{s}^{-1}$ to 0.095 $\text{nm}\,\text{s}^{-1}$, as $c_{\text{dop},\text{b}}(\infty)$ is raised from $4\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$.
@@
In the sintering process, the driving forces are initially large and decrease as the grain size increases. In other words, the GB core velocity starts on the upper right side of the S-shaped curves in Fig. 5 at the initial stage of the grain growth. As the grain size increases, the GB core velocity decreases to the critical velocity. At this point, a transition from high velocity to low velocity occurs at the critical value. In Fig. 5(a), we observe GB core velocity jumps when $\Delta g_{\text{V}_{\text{O}}^{..}}=$ -1.5 eV and -1.75 eV. The critical velocities for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ are approximately 0.35 $\text{nm}\,\text{s}^{-1}$. Additionally, the GB core velocity jumps to a lower value when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is more negative. In Fig. 5(b), the critical GB core velocity increases from 0.15 $\text{nm}\,\text{s}^{-1}$ to 0.65 $\text{nm}\,\text{s}^{-1}$ as $\Delta g_{\text{dop}}$ decreases from -0.3 eV to -0.7 eV. In Fig. 5(c), a larger $c_{\text{dop},\text{b}}(\infty)$ leads to a higher critical velocity.
In summary, the formation of SCLs at the quasi-equilibrium state has been systematically investigated under varying GB velocities and defect chemistry parameters. We conclude that (i) higher GB velocities lead to reduced dopant segregation in the GB core, more pronounced asymmetry in the oxygen vacancy concentration profiles, increased GB potential, and an elongated SCL length in the growing grain. (ii) the segregation energy of the dopant influences the concentration distribution pattern within the SCL and the GB potential, particularly at intermediate GB velocities. However, at extremely high GB velocities, its impact becomes negligible. (iii) the fundamental relationship between defect chemistry parameters and solute drag effects has been elucidated. When $\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$ is smaller than 0.5 eV, the solute drag effect is negligible. However, at low $\lvert\Delta g_{\text{dop}}\rvert$ and dopant concentrations, minimal solute drag effects can still be observed. Thus, reducing the segregation energy of oxygen vacancy emerges as an effective strategy to minimize solute drag effects.
@@
5.3 Polycrystalline simulations during sintering